OpenStudy (anonymous):

A rock thrown into a pond causes a circular ripple. if the radius of the ripple is increasing by 3ft/sec how fast is the area changing when the radius is 12 feet?

OpenStudy (anonymous):

Area of a circle: \[ A = \pi r^2 \]Can you differentiate with respect to time?

OpenStudy (anonymous):

Consider that area and radius are functions of time: \[ A(t) = \pi [r(t)]^2 \]

OpenStudy (anonymous):

the derivative would be 2πrr'(t) ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[ A'(t) = 2\pi r(t) r'(t) \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

is the last equation to solve it 6π(12)?

OpenStudy (anonymous):

Yeah, you can simplify it more, but that is correct.

OpenStudy (anonymous):

so the answer would be 226.08ft/sec

OpenStudy (anonymous):

Answer is \(72\pi\; \text{ft}^2/\text{s}\)

OpenStudy (anonymous):

wio im lost from A′(t)=2πr(t)r′(t) what numbers go in there

OpenStudy (anonymous):

\(r'(t) = 3 \;\text{ft/s}\) and \(r(t) = 12\; \text{ft}\)

OpenStudy (anonymous):

okay so it would be 2π(12)(3)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how did u get 72 for the answer